OpenStudy (anonymous):

Please Help Will Medal And Fan! For standard position angle determined by the point (x,y) what are the values of the trigonometric functions. For the point (5,12) find csc theta and sec theta

OpenStudy (anonymous):

A. Csc theta= 13/12 , sec theta 12/13 B. Csc theta= 12/13, sec theta=13/5 C. Csc theta= 13/12, sec theta=15/5 D. Csc theta= 5/12, sec theta=13/5

OpenStudy (anonymous):

a = 9, b = 12, c = 15 and, If "theta" is the angle between a & b, then cos(theta) = 15/12 = 5/4 and, sin(theta) = 15/9 = 5/3 So, sec(theta) = 1/cos(theta) = 1/(5/4) = 4/5 csc(theta) = 1/sin(theta) = 1/(5/3) = 3/5

OpenStudy (anonymous):

@Johnica Those are none of my answer choices

OpenStudy (anonymous):

?

OpenStudy (anonymous):

csc = 9/15 sec = 12/15

OpenStudy (anonymous):

i am confused it still doesn't match my any of my answer choice is there something else i have to do?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok is it D?

OpenStudy (anonymous):

@Johnica

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

someone help me?@anyone

OpenStudy (anonymous):

@izzy529

OpenStudy (anonymous):

fine